# Pruning the Paradox: How CLIP’s Most Informative Heads Enhance Performance While Amplifying Bias

### v2

@@ Abstract
CLIP is one of the most popular foundational models and is heavily used for many vision-language tasks. However, little is known about the inner workings of CLIP. While recent work has proposed decomposition-based interpretability methods for identifying textual descriptions of attention heads in CLIP, the implications of conceptual consistency in these text labels on interpretability and model performance has not been explored. To bridge this gap, we study the conceptual consistency of text descriptions for attention heads in CLIP-like models. We conduct extensive experiments on six different models from OpenAI and OpenCLIP which vary by size, type of pre-training data and patch size. We propose Concept Consistency Score (CCS), a novel interpretability metric that measures how consistently individual attention heads in CLIP models align with specific concepts. To assign concept labels to heads, we use in-context learning with ChatGPT, guided by a few manually-curated examples, and validate these labels using an LLM-as-a-judge approach. Our soft-pruning experiments reveal that high CCS heads are critical for preserving model performance, as pruning them leads to a significantly larger performance drop than pruning random or low CCS heads. Notably, we find that high CCS heads capture essential concepts and play a key role in out-of-domain detection, concept-specific reasoning, and video-language understanding. Moreover, we prove that high CCS heads learn spurious correlations amplifying social biases. These results position CCS as a powerful interpretability metric exposing the paradox of performance and social biases in CLIP models.
Machine Learning, ICML
@@ 1 Introduction
Large-scale vision-language (VL) models such as CLIP  (Radford et al., 2021) have significantly advanced state-of-the-art performance in vision tasks in recent years. Consequently, CLIP has been extensively used as a foundational model for downstream tasks such as video retrieval, image generation, and segmentation  (Luo et al., 2022; Liu et al., 2024; Brooks et al., 2023; Esser et al., 2024; Kirillov et al., 2023). This has enabled the construction of compositional models combining CLIP with other foundation models, thereby increasing the functionality of CLIP while also adding complexity to the overall model structure. However, as these models gain prominence in real-world applications, their embedded social biases  (Howard et al., 2024; Hall et al., 2023; Seth et al., 2023) have emerged as a critical concern with potentially harmful downstream consequences. Despite the growing body of work documenting these biases, a fundamental question remains unanswered: what mechanisms within these models’ architectures drive both their impressive capabilities and problematic shortcomings?
Recent interpretability advances  (Gandelsman et al., ) have made initial progress by decomposing CLIP’s image representations into contributions from individual attention heads, identifying text sequences that characterize different heads’ semantic roles. However, this approach provides only a partial view into CLIP’s inner workings, leaving a critical missing piece: systematic understanding of the visual concepts encoded at the attention head level—and how these concepts underpin both the model’s strengths and its social failures.
Our work addresses this critical gap through a novel interpretability framework we call “conceptual consistency”. This framework systematically analyzes which visual concepts are learned by individual attention heads and how consistently these concepts are processed throughout the model’s architecture. First, we identify interpretable structures within the individual heads of the last four layers of the model using a set of text descriptions. To accomplish this, we employ the TextSpan algorithm  (Gandelsman et al., ), which helps us find the most appropriate text descriptions for each head. After identifying these text descriptions, we assign labels to each head representing the common property shared by the descriptions. This labeling process is carried out using in-context learning with ChatGPT. We begin by manually labeling five pairs of text descriptions and their corresponding concept labels, which serve as examples. These examples are then used to prompt ChatGPT to assign labels for the remaining heads.
@@
Figure 1: Figure shows the steps of computing Concept Consistency Score for each head.
@@ 2 Related Work
Early research on interpretability primarily concentrated on convolutional neural networks (CNNs) due to their intricate and opaque decision-making processes (Zeiler & Fergus, 2014; Selvaraju et al., 2017; Simonyan et al., 2014; Fong & Vedaldi, 2017; Hendricks et al., 2016). More recently, the interpretability of Vision Transformers (ViT) has garnered significant attention as these models, unlike CNNs, rely on self-attention mechanisms rather than convolutions. Researchers have focused on task-specific analyses in areas such as image classification, captioning, and object detection to understand how ViTs process and interpret visual information (Dong et al., 2022; Elguendouze et al., 2023; Mannix & Bondell, 2024; Xue et al., 2022; Cornia et al., 2022; Dravid et al., 2023). One of the key metrics used to measure interpretability in ViTs is the attention mechanism itself, which provides insights into how the model distributes focus across different parts of an image when making decisions (Cordonnier et al., 2019; Chefer et al., 2021). This has led to the development of techniques that leverage attention maps to explain ViT predictions. Early work on multimodal interpretability, which involves models that handle both visual and textual inputs, probed tasks such as how different modalities influence model performance (Cao et al., 2020; Madasu & Lal, 2023) and how visual semantics are represented within the model (Hendricks & Nematzadeh, 2021; Lindström et al., 2021). Aflalo et al. (Aflalo et al., 2022) explored interpretability methods for vision-language transformers, examining how these models combine visual and textual information to make joint decisions. Similarly, Stan et al. (Stan et al., 2024) proposed new approaches for interpreting vision-language models, focusing on the interactions between modalities and how these influence model predictions. Our work builds upon and leverages the methods introduced by Gandelsman et al. ; Gandelsman et al. (2024) to interpret attention heads, neurons, and layers in vision-language models, providing deeper insights into their decision-making processes.
@@ 3.1 Preliminaries
In this section, we describe our methodology, starting with the TextSpan  (Gandelsman et al., ) algorithm and its extension across all attention heads in multiple CLIP models using in-context learning. TextSpan associates each attention head with relevant text descriptions by analyzing the variance in projections between head outputs and candidate text representations. Through iterative projections, it identifies distinct components aligned with different semantic aspects. While effective at linking heads to descriptive text spans, TextSpan does not assign explicit concept labels. In the next section, we detail our method for labeling the concepts learned by individual CLIP heads.
@@ 3.2 Concept Consistency Score (CCS)
We introduce the Concept Consistency Score (CCS) as a systematic metric for analyzing the concepts (properties) learned by transformer layers and attention heads in CLIP-like models. This score quantifies the alignment between the textual representations produced by a given head and an assigned concept label. Figure  1 illustrates our approach, with the following sections detailing each step in computing CCS.
@@ 3.2.3 Evaluating Concept Consistency
where $\mathbb{1}[\cdot]$ is an indicator function that returns 1 if $T_{i}$ to be consistent with $C_{h}$, and 0 otherwise. To ensure a high standard of reliability, we define consistency strictly—only if all three LLM judges independently rate $T_{i}$ as consistent with $C_{h}$. This requirement for unanimous agreement minimizes the influence of individual model biases or variability in judgment  (Liu & Zhang, 2025), thereby enhancing the robustness and trustworthiness of the overall concept consistency score.
@@
Next, we categorize each attention head based on its Concept Consistency Score (CCS) into three levels: high, moderate, and low. A head is considered to have a high CCS if all of its associated text descriptions align with the labeled concept, indicating that the head is highly specialized and likely encodes features relevant to that concept. Moderate CCS heads exhibit partial alignment, with three out of five text descriptions matching the concept label, suggesting that they capture the concept to some extent but not exclusively. In contrast, low CCS heads have zero or only one matching description, implying minimal relevance and indicating that these heads are largely unrelated to the given concept. This categorization provides insight into the degree of concept selectivity exhibited by individual attention heads. Table  1 shows examples of different types of CCS heads.
@@ 3.4.1 Tasks
Image classification: CIFAR-10  (Krizhevsky et al., 2009), CIFAR-100  (Krizhevsky et al., 2009), Food-101  (Bossard et al., 2014), Country-211  (Radford et al., 2021) and Oxford-pets  (Parkhi et al., 2012).
Out-of-domain classification: Imagenet-A  (Hendrycks et al., 2021b) and Imagenet-R  (Hendrycks et al., 2021a).
Video retrieval: MSRVTT  (Xu et al., 2016), MSVD  (Chen & Dolan, 2011), DiDeMo  (Anne Hendricks et al., 2017).
Bias: FairFace  (Karkkainen & Joo, 2021), SocialCounterFactuals  (Howard et al., 2024).
@@ 3.4.2 Models
For experiments we use the following six foundational image-text models: ViT-B-32, ViT-B-16 and ViT-L-14 pretrained from OpenAI-400M  (Radford et al., 2021) and LAION2B  (Schuhmann et al., 2022). Next, we discuss in detail the results from the experiments.
@@ 4 Results and Discussion
In this section we examine the role of the Concept Consistency Score (CCS) in revealing CLIP’s decision-making process, focusing on the question: How does CCS provide deeper insights into the functional role of individual attention heads in influencing downstream tasks? To explore this, we perform a soft-pruning analysis by zeroing out attention weights of heads with extreme CCS values—specifically, high CCS (CCS = 5) and low CCS (CCS $\leq$ 1). This approach disables selected heads without modifying the model architecture. As shown in Table  3, pruning high-CCS heads consistently causes significant drops in zero-shot classification performance across CIFAR-10, CIFAR-100 and FOOD-101 while pruning low-CCS heads has a minimal effect. This performance gap demonstrates that CCS effectively identifies heads encoding critical, concept-aligned information, making it a reliable tool for interpreting CLIP’s internal decision-making mechanisms.
@@
4.2 High CCS vs random heads pruning
In the previous section, we showed that attention heads with high Concept Consistency Scores (CCS) are crucial to CLIP’s performance. To validate whether these heads are truly more important than others, we perform a controlled comparison against random pruning. Specifically, we randomly prune the same number of attention heads—excluding high-CCS heads—and repeat this across five seeds, averaging the results. As illustrated in Figure  2, pruning high-CCS heads consistently causes a significantly larger drop in zero-shot accuracy compared to random pruning across datasets and model variants. In contrast, random pruning results in only minor performance degradation, highlighting the functional importance of high-CCS heads. Interestingly, we also find that larger CLIP models show a smaller performance gap between high-CCS and random pruning, suggesting that larger architectures may be more robust due to greater redundancy or more distributed representations. These findings support CCS as a reliable and interpretable metric for identifying concept-relevant heads and offer deeper insights into how CLIP organizes conceptual information.
@@ 4.3 High CCS heads are crucial for out-of-domain (OOD) detection
While our earlier experiments primarily focused on in-domain datasets such as CIFAR-10 and CIFAR-100 to validate the Concept Consistency Score (CCS), understanding model behavior under out-of-domain (OOD) conditions is a critical step toward evaluating models’ robustness. Table  4 demonstrates the results on ImageNet-A and ImageNet-R datasets respectively. From the table, we observe that pruning heads with high CCS scores leads to a substantial degradation in model performance, underscoring the critical role these heads play in the model’s decision-making process. Notably, the ViT-B-16-OpenAI model exhibits the most pronounced drop in performance upon pruning high CCS heads, suggesting that this model relies heavily on a smaller set of concept-specific heads for robust feature representation consistent with the observations previously. These results demonstrate that CCS is a powerful metric for identifying attention heads that encode essential, generalizable concepts in CLIP models.
@@
To investigate the functional role of high Concept Consistency Score (CCS) heads, we conduct concept-specific pruning experiments. In these experiments, we prune heads with high CCS scores corresponding to a target concept (e.g., locations) and evaluate the model’s performance on tasks aligned with that concept, such as location classification. In contrast, we also prune heads associated with unrelated concepts (e.g., animals) and assess the resulting impact on task performance. Our results indicate that pruning high CCS heads leads to a significant drop in task performance, validating that these heads encode essential concept-relevant information. For instance, in the ViT-B-16 model, pruning location heads results in a substantial decrease in location classification accuracy from 22.81% to 14.09%, as shown in Figure  3. Conversely, pruning heads corresponding to unrelated concepts has little effect on performance, demonstrating the concept-specific nature of high CCS heads, as illustrated in Figure  4.
@@
To further assess the importance of high CCS heads for downstream tasks, we conducted a series of zero-shot video retrieval experiments on three popular datasets: MSRVTT, MSVD, and DIDEMO under different pruning strategies. Figure  5 shows the results of this experiment. Notably, pruning high CCS (Concept Consistency Score) heads consistently leads to a substantial drop in performance across all datasets, demonstrating their critical role in preserving CLIP’s retrieval capabilities. For instance, on MSRVTT and MSVD, high CCS pruning significantly underperforms compared to low CCS and random head pruning, which show much milder performance degradation. Interestingly, low CCS and random head pruning maintain performance much closer to the original unpruned model, indicating that not all attention heads contribute equally to model competence. This consistent trend across datasets highlights that heads with high CCS scores are essential for encoding concept-aligned information necessary for accurate zero-shot video retrieval.
Occupations
Prompt Example
A <occupation> A biologist
A photo of <occupation> A photo of biologist
A picture of <occupation> A picture of biologist
An image of <occupation> An image of biologist
Model Race Gender
Original High CCS Original High CCS
ViT-B-32-OpenAI 61.75 60.47 41.24 41.11
ViT-B-32-datacomp 49.2 48.32 21.35 20.97
ViT-B-16-OpenAI 64.61 55.55 40.19 38.21
ViT-B-16-LAION 59.21 56.72 43.55 43.11
ViT-L-14-OpenAI 59.28 59.75 34.7 32.23
ViT-L-14-LAION 61.92 55.59 43.02 39.17
@@
Model Race Gender
Original High CCS Original High CCS
ViT-B-32-OpenAI 3.65 2.43 4.05 1.22
ViT-B-16-OpenAI 2.43 1.22 0.81 2.03
ViT-L-14-OpenAI 2.03 0.81 2.42 1.62

### v3

@@ Abstract
CLIP is one of the most popular foundation models and is heavily used for many vision-language tasks, yet little is known about its inner workings. As CLIP is increasingly deployed in real-world applications, it is becoming even more critical to understand its limitations and embedded social biases to mitigate potentially harmful downstream consequences. However, the question of what internal mechanisms drive both the impressive capabilities as well as problematic shortcomings of CLIP has largely remained unanswered. To bridge this gap, we study the conceptual consistency of text descriptions for attention heads in CLIP-like models. Specifically, we propose Concept Consistency Score (CCS), a novel interpretability metric that measures how consistently individual attention heads in CLIP models align with specific concepts. Our soft-pruning experiments reveal that high CCS heads are critical for preserving model performance, as pruning them leads to a significantly larger performance drop than pruning random or low CCS heads. Notably, we find that high CCS heads capture essential concepts and play a key role in out-of-domain detection, concept-specific reasoning, and video-language understanding. Moreover, we prove that high CCS heads learn spurious correlations which amplify social biases. These results position CCS as a powerful interpretability metric exposing the paradox of performance and social biases in CLIP models.
Pruning the Paradox: How CLIP’s Most Informative Heads Enhance Performance While Amplifying Bias
Avinash Madasu   Vasudev Lalthanks: Work partially completed while at Intel Labs.    Phillip Howard11footnotemark: 1 Intel Labs   Oracle   Thoughtworks   avinash.madasu@intel.com   phillip.howard@thoughtworks.com
@@ 1 Introduction
Large-scale vision-language (VL) models such as CLIP  Radford et al. (2021) have significantly advanced state-of-the-art performance in vision tasks in recent years. Consequently, CLIP has been extensively used as a foundational model for downstream tasks such as video retrieval, image generation, and segmentation  Luo et al. (2022); Liu et al. (2024); Brooks et al. (2023); Esser et al. (2024); Kirillov et al. (2023). This has enabled the construction of compositional models combining CLIP with other foundation models, thereby increasing the functionality of CLIP while also adding complexity to the overall model structure. However, as these models gain prominence in real-world applications, their embedded social biases  Howard et al. (2024); Hall et al. (2023); Seth et al. (2023) have emerged as a critical concern with potentially harmful consequences. Despite the growing body of work documenting these biases, a fundamental question remains: what mechanisms within these models’ architectures drive both their impressive capabilities and problematic shortcomings?
Recent interpretability advances  Gandelsman et al. have made initial progress by decomposing CLIP’s image representations into contributions from individual attention heads, identifying text sequences that characterize different heads’ semantic roles. However, this approach provides only a partial view into CLIP’s inner workings, leaving a critical missing piece: systematic understanding of the visual concepts encoded at the attention head level—and how these concepts underpin both the model’s strengths and its social failures.
Our work addresses this critical gap through a novel interpretability framework we call “conceptual consistency”. This framework systematically analyzes which visual concepts are learned by individual attention heads and how consistently these concepts are processed throughout the model’s architecture. First, we identify interpretable structures within the individual heads of the last four layers of the model using a set of text descriptions. To accomplish this, we employ the TextSpan algorithm  Gandelsman et al. , which helps us find the most appropriate text descriptions for each head. After identifying these text descriptions, we assign labels to each head representing the common property shared by the descriptions. This labeling process is carried out using in-context learning with ChatGPT. We begin by manually labeling five pairs of text descriptions and their corresponding concept labels, which serve as examples. These examples are then used to prompt ChatGPT to assign labels for the remaining heads.
@@
Figure 1: Illustration of our approach to computing Concept Consistency Score for each attention head.
@@
Model High Medium Low
ViT-B-32-OpenAI 8 13 4
ViT-B-32-datacomp 11 9 8
ViT-B-16-OpenAI 10 14 4
ViT-B-16-LAION 13 12 6
ViT-L-14-OpenAI 16 13 11
ViT-L-14-LAION 21 14 3
Table 2: Count of high, medium and low CCS heads in CLIP models.
Model CIFAR-10 CIFAR-100
Original K=5 K=7 K=9 K=11 K=13 Original K=5 K=7 K=9 K=11 K=13
ViT-B-32-OpenAI 75.68 71.31 71.86 73.66 74.53 70.54 65.08 56.07 62.21 60.17 56.42 59.05
ViT-B-32-datacomp 72.07 70.50 70.59 69.61 69.93 70.20 54.95 53.14 53.59 53.73 53.81 53.91
ViT-B-16-OpenAI 78.10 63.93 71.83 76.08 64.67 76.74 68.22 51.70 52.45 61.90 62.59 55.54
ViT-B-16-LAION 82.82 78.91 81.25 80.25 78.03 78.56 76.92 65.55 69.94 72.46 67.59 69.49
ViT-L-14-OpenAI 86.94 86.29 85.44 85.89 85.49 85.58 78.28 75.66 74.85 74.70 76.73 77.19
ViT-L-14-LAION 88.29 86.48 85.99 86.34 87.28 85.99 83.37 80.07 80.75 82.03 82.14 80.11
Table 3: Accuracy (%) on CIFAR-10 and CIFAR-100 datasets across different values of TEXTSPAN ($K$).
@@ 2 Related Work
Early research on interpretability primarily concentrated on convolutional neural networks (CNNs) due to their intricate and opaque decision-making processes (Zeiler and Fergus, 2014; Selvaraju et al., 2017; Simonyan et al., 2014; Fong and Vedaldi, 2017; Hendricks et al., 2016). More recently, the interpretability of Vision Transformers (ViT) has garnered significant attention as these models, unlike CNNs, rely on self-attention mechanisms rather than convolutions. Researchers have focused on task-specific analyses in areas such as image classification, captioning, and object detection to understand how ViTs process and interpret visual information Dong et al. (2022); Elguendouze et al. (2023); Mannix and Bondell (2024); Xue et al. (2022); Cornia et al. (2022); Dravid et al. (2023). One of the key metrics used to measure interpretability in ViTs is the attention mechanism itself, which provides insights into how the model distributes focus across different parts of an image when making decisions Cordonnier et al. (2019); Chefer et al. (2021). This has led to the development of techniques that leverage attention maps to explain ViT predictions. Early work on multimodal interpretability, which involves models that handle both visual and textual inputs, probed tasks such as how different modalities influence model performance Cao et al. (2020); Madasu and Lal (2023) and how visual semantics are represented within the model Hendricks and Nematzadeh (2021); Lindström et al. (2021). Aflalo et al. (Aflalo et al., 2022) explored interpretability methods for vision-language transformers, examining how these models combine visual and textual information to make joint decisions. Similarly, Stan et al. (Stan et al., 2024) proposed new approaches for interpreting vision-language models, focusing on the interactions between modalities and how these influence model predictions. Our work builds upon and leverages the methods introduced by Gandelsman et al. ; Gandelsman et al. (2024) to interpret attention heads, neurons, and layers in vision-language models, providing deeper insights into their decision-making processes. Several works  Shu et al. (2023); Mayilvahanan et al. ; Moeller et al. (2024) explored the out-of-domain generalization of CLIP models on several benchmarks. In contrast to these works, our work provides a new interpretability metric to understand the decision making processes of attention heads in CLIP models.
@@ 3.1 Preliminaries
In this section, we describe our methodology, starting with the TextSpan  Gandelsman et al. algorithm and its extension across all attention heads in multiple CLIP models using in-context learning. Specifically, we deal with attention heads in the image encoder. TextSpan associates each attention head with relevant text descriptions by analyzing the variance in projections between head outputs and candidate text representations. Through iterative projections, it identifies distinct components aligned with different semantic aspects. While effective at linking heads to descriptive text spans, TextSpan does not assign explicit concept labels. We inherited the set of 3498 text spans from Gandelsman et al.  Gandelsman et al. , which were generated by prompting ChatGPT-3.5 to produce general image descriptions using the prompts shown below. After obtaining an initial set, ChatGPT was manually prompted to generate more examples of specific patterns we found (e.g., image with a lunar eclipse, Oceanic coral reef). In the next section, we detail our method for labeling the concepts learned by individual CLIP heads.
@@ 3.2 Concept Consistency Score (CCS)
We introduce the Concept Consistency Score (CCS) as a systematic metric for analyzing the concepts (properties) learned by transformer layers and attention heads in CLIP-like models. This score quantifies the alignment between the textual representations produced by a given head and an assigned concept label. Existing interpretability methods  Chefer et al. (2021); Abnar and Zuidema (2020); Barkan et al. (2021) are largely descriptive and do not identify which internal components (e.g., attention heads) are reliably responsible for model behavior. This lack of discriminative insight limits their utility in scenarios where intervention or control is required. CCS addresses this by measuring concept consistency across heads, enabling us to isolate specialized heads that consistently encode particular concepts. This makes interpretability actionable by supporting tasks like pruning, debugging, or understanding failure modes with certainty and control. Figure  1 illustrates our approach, with the following sections detailing each step in computing CCS.
@@ 3.2.3 Evaluating Concept Consistency
where $\mathbb{1}[\cdot]$ is an indicator function that returns 1 if $T_{i}$ to be consistent with $C_{h}$, and 0 otherwise. To ensure a high standard of reliability, we define consistency strictly—only if all three LLM judges independently rate $T_{i}$ as consistent with $C_{h}$. This requirement for unanimous agreement minimizes the influence of individual model biases or variability in judgment  Liu and Zhang (2025), thereby enhancing the robustness and trustworthiness of the overall concept consistency score.
@@
Next, we categorize each attention head based on its Concept Consistency Score (CCS) into three levels: high, moderate, and low. A head is considered to have a high CCS if all of its associated text descriptions align with the labeled concept, indicating that the head is highly specialized and likely encodes features relevant to that concept. Moderate CCS heads exhibit partial alignment, with three out of five text descriptions matching the concept label, suggesting that they capture the concept to some extent but not exclusively. In contrast, low CCS heads have zero or only one matching description, implying minimal relevance and indicating that these heads are largely unrelated to the given concept. This categorization provides insight into the degree of concept selectivity exhibited by individual attention heads. Table  1 shows examples of different types of CCS heads and Table  2 shows the count of high, moderate and low CCS heads in all CLIP models.
@@ 3.3 Evaluating LLM Judgment Alignment with Human Annotations
3.4 Sensitivity of CCS to the number of TextSpan descriptions
Next, we analyze the robustness of CCS computation to number of TextSpan ($K$) descriptions. To directly address this issue, we conduct experiments evaluating CCS performance across different numbers of TextSpan descriptions on CIFAR-10 and CIFAR-100 datasets. The results are depicted in the table  3. From the results, it is evident that performance remains remarkably stable across all $K$ values, with variations typically within 1-3% accuracy points across multiple CLIP models. Importantly, the relative performance ordering between different model configurations remains consistent regardless of $K$, indicating that our concept alignment assessment captures meaningful patterns. Performance does not consistently improve with higher $K$ values, suggesting that five TextSpan descriptions capture sufficient conceptual diversity without unnecessary computational overhead. These findings directly validate that CCS is robust to the number of TextSpan descriptions and not overly sensitive to this hyperparameter choice.
@@ 3.5.1 Tasks
For our experiments, we use a wide variety of datasets focusing on the tasks of image classification, out-of-domain classification, video retrieval, and bias measurement. Below, we mention datasets for each of the tasks.
Image classification: CIFAR-10  Krizhevsky et al. (2009), CIFAR-100  Krizhevsky et al. (2009), Food-101  Bossard et al. (2014), Country-211  Radford et al. (2021) and Oxford-pets  Parkhi et al. (2012).
Out-of-domain classification: Imagenet-A  Hendrycks et al. (2021b) and Imagenet-R  Hendrycks et al. (2021a).
Video retrieval: MSRVTT  Xu et al. (2016), MSVD  Chen and Dolan (2011), DiDeMo  Anne Hendricks et al. (2017).
Bias: FairFace  Karkkainen and Joo (2021), SocialCounterFactuals  Howard et al. (2024).
@@ 3.5.2 Models
For experiments we use the following six foundational image-text models: ViT-B-32, ViT-B-16 and ViT-L-14 pretrained from OpenAI-400M  Radford et al. (2021) and LAION2B  Schuhmann et al. (2022). Next, we discuss in detail the results from the experiments.
@@
Model CIFAR-100 FOOD-101 Country-211
ViT-B-32-OpenAI 65.08 62.39 58.58 84.01 82.12 73.24 17.16 16.3 12.23
ViT-B-32-datacomp 54.95 53.72 53.77 41.66 40.77 39.31 4.43 4.37 4.41
ViT-B-16-OpenAI 68.22 65.38 56.17 88.73 87.36 83.23 22.81 21.79 21.74
ViT-B-16-LAION 76.92 72.51 72.73 86.63 81.4 80.51 20.45 16.87 14.39
ViT-L-14-OpenAI 78.28 77.55 73 93.07 92.79 90.15 31.91 30.63 25.38
ViT-L-14-LAION 83.37 83.25 83.35 91.02 90.35 90.25 26.41 25.66 25.43
Table 7: Accuracy comparison of various CLIP models on CIFAR-100, FOOD-101 and Country-211 datasets. The values represent original accuracy, performance after pruning equal number of high-CCS and low-CCS heads.
@@ 4 Results and Discussion
In this section we examine the role of the Concept Consistency Score (CCS) in revealing CLIP’s decision-making process, focusing on the question: How does CCS provide deeper insights into the functional role of individual attention heads in influencing downstream tasks? To explore this, we perform a soft-pruning analysis by zeroing out attention weights of heads with extreme CCS values—specifically, high CCS (CCS = 5) and low CCS (CCS $\leq$ 1). This approach disables selected heads without modifying the model architecture. As shown in Table  5, pruning high-CCS heads consistently causes significant drops in zero-shot classification performance across CIFAR-10, CIFAR-100 and FOOD-101 while pruning low-CCS heads has a minimal effect. This performance gap demonstrates that CCS effectively identifies heads encoding critical, concept-aligned information, making it a reliable tool for interpreting CLIP’s internal decision-making mechanisms.
@@
4.2 Pruning equal number of high CCS, low CCS and random heads
In the previous section, we showed that attention heads with high Concept Consistency Scores (CCS) are crucial to CLIP’s performance. To validate whether these heads are truly more important than others, we perform a controlled comparison against random pruning. Specifically, we randomly prune the same number of attention heads—excluding high-CCS heads—and repeat this across five seeds, averaging the results. As illustrated in Figure  2, pruning high-CCS heads consistently causes a significantly larger drop in zero-shot accuracy compared to random pruning across datasets and model variants. In contrast, random pruning results in only minor performance degradation, highlighting the functional importance of high-CCS heads. Interestingly, we also find that larger CLIP models show a smaller performance gap between high-CCS and random pruning, suggesting that larger architectures may be more robust due to greater redundancy or more distributed representations.
Similarly, we conducted experiments where we pruned an equal number of high and low CCS attention heads across multiple datasets (CIFAR-100, FOOD-101, Country-211). Results are shown in the table  7. From the table, we observe that pruning high-CCS heads leads to a substantially larger performance drop, even when the number of pruned heads is held constant. This effectively rules out the explanation that the observed degradation is merely due to pruning more heads in the high-CCS condition. Taken together, these findings support CCS as a reliable and interpretable metric for identifying concept-relevant heads and offer deeper insights into how CLIP organizes conceptual information.
@@ 4.3 High CCS heads are crucial for out-of-domain (OOD) detection
While our earlier experiments primarily focused on in-domain datasets such as CIFAR-10 and CIFAR-100 to validate the Concept Consistency Score (CCS), understanding model behavior under out-of-domain (OOD) conditions is a critical step toward evaluating models’ robustness. Table  6 demonstrates the results on ImageNet-A and ImageNet-R datasets respectively. From the table, we observe that pruning heads with high CCS scores leads to a substantial degradation in model performance, underscoring the critical role these heads play in the model’s decision-making process. Notably, the ViT-B-16-OpenAI model exhibits the most pronounced drop in performance upon pruning high CCS heads, suggesting that this model relies heavily on a smaller set of concept-specific heads for robust feature representation consistent with the observations previously. These results demonstrate that CCS is a powerful metric for identifying attention heads that encode essential, generalizable concepts in CLIP models.
@@
To investigate the functional role of high Concept Consistency Score (CCS) heads, we conduct concept-specific pruning experiments. In these experiments, we prune heads with high CCS scores corresponding to a target concept (e.g., locations) and evaluate the model’s performance on tasks aligned with that concept, such as location classification. In contrast, we also prune heads associated with unrelated concepts (e.g., animals) and assess the resulting impact on task performance. Our results indicate that pruning high CCS heads leads to a significant drop in task performance, validating that these heads encode essential concept-relevant information. For instance, in the ViT-B-16 model, pruning location heads results in a substantial decrease in location classification accuracy from 22.81% to 14.09%, as shown in Figure 3. Conversely, pruning heads corresponding to unrelated concepts has little effect on performance, demonstrating the concept-specific nature of high CCS heads, as illustrated in Figure 4.
@@
To further assess the importance of high CCS heads for downstream tasks, we conducted zero-shot video retrieval experiments on three popular datasets (MSRVTT, MSVD, and DIDEMO) under different pruning strategies. The results in Figure  5 show that pruning high CCS (Concept Consistency Score) heads consistently leads to a substantial drop in performance across all datasets, demonstrating their critical role in preserving CLIP’s retrieval capabilities. For instance, on MSRVTT and MSVD, high CCS pruning significantly underperforms compared to low CCS and random head pruning, which show much milder performance degradation. Interestingly, low CCS and random head pruning maintain performance much closer to the original unpruned model, indicating that not all attention heads contribute equally to model competence. This consistent trend across datasets highlights that heads with high CCS scores are essential for encoding concept-aligned information necessary for accurate zero-shot video retrieval.
Model Race ($\downarrow$) Gender ($\downarrow$)
Original High CCS Low CCS Original High CCS Low CCS
ViT-B-32-OpenAI 61.8 60.5 65.0 41.2 41.1 42.5
ViT-B-32-datacomp 49.2 48.3 49.7 21.7 21 21.6
ViT-B-16-OpenAI 64.6 55.6 66.6 40.2 38.2 40.5
ViT-B-16-LAION 59.2 56.7 56.6 43.6 43.1 43.8
ViT-L-14-OpenAI 59.3 59.8 59.9 34.7 32.2 35.2
ViT-L-14-LAION 61.9 55.6 59.7 43 39.2 43.9
@@
Model Race ($\downarrow$) Gender ($\downarrow$)
Original High CCS Low CCS Original High CCS Low CCS
ViT-B-32-OpenAI 4.1 1.2 2.4 3.7 2.4 1.2
ViT-B-16-OpenAI 0.8 2.0 1.2 2.4 1.2 2.1
ViT-L-14-OpenAI 2.0 0.8 2.0 2.4 2.0 1.6